OpenStudy (anonymous):

Choose one of the factors of 27x^3 + 512y^3. 3 3x - 8y 9x^2 - 24xy + 64y^2 9x2 + 24xy + 64y2

OpenStudy (anonymous):

Also, I will give a medal, and if you could teach me how to do it I would love it!

OpenStudy (anonymous):

I am pretty sure it is 3x - 8y you will want to check to make sure though

OpenStudy (anonymous):

Can you tell me how to get that answer?

OpenStudy (anonymous):

let me look more into it and I will get back to you

OpenStudy (anonymous):

use the formula : a^3 + b^3 = (a+b)(a^2 - ab + b^2) 27x^3 + 512y^3 = (3x)^3 + (8y)^3 now, use the formula above to get of factors a^3 + b^3 = (a+b)(a^2 - ab + b^2) we have (3x)^3 + (8y)^3 it means a = 3x and b = 8y therefore, (3x)^3 + (8y)^3 = (3x+8y)((3x)^2 - 3x(8y) + (8y)^2) (3x)^3 + (8y)^3 = (3x+8y)(9x^2 - 24xy + 64y^2) so, it is C I found your problem already answered hopefully this helps.